OpenStudy (anonymous):

what is the simplest form of √128x^5y^2

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

split 128 in such a way that you can get a complete square like in last problem 9 was square of 3 thats why 63=9*7

OpenStudy (anonymous):

you can do this

OpenStudy (anonymous):

4 and 8

OpenStudy (anonymous):

\[\sqrt{128x^{5}y^2}\] 128 can be divided as 2*64 or \[2*8^{2}\] and x^5 as \[(x^2)^2*x\]

OpenStudy (anonymous):

So\[\sqrt{2*8^2*(x^2)^2*x*y^2}\]

OpenStudy (anonymous):

^^^ thats the answer ? im so confused

OpenStudy (anonymous):

4 and 8? no see sandeep solved it there

OpenStudy (anonymous):

no thats not the answer

OpenStudy (anonymous):

you take complete square terms outside like 64 = 8^2 when you take it out of root you get 8

OpenStudy (anonymous):

when you take y^2 out of root you get y

OpenStudy (anonymous):

and when you take x^4 out of root you get x^2

OpenStudy (anonymous):

we have a relation \[\sqrt{a*b*c} = \sqrt{a}*\sqrt{b}*\sqrt{c}\]

OpenStudy (anonymous):

Im so confused/:

OpenStudy (anonymous):

\[\sqrt{2*8^2*(x^2)^2*x*y^2}\] = \[\sqrt{2}*\sqrt{8^2}*\sqrt{(x^2)^2}*\sqrt{x}*\sqrt{y^2}\]

OpenStudy (anonymous):

that gives u \[\sqrt{2}*8*x^2*\sqrt{x}*y\]

OpenStudy (anonymous):

\[8\sqrt{2}x^{3/2}y\]

OpenStudy (anonymous):

@miah did u get it?

OpenStudy (anonymous):

hmm could hv been easier to teach face to face answer is \[8x^2y \sqrt{2xy}\]

OpenStudy (anonymous):

Yes i too feel same @aajugdar

OpenStudy (anonymous):

Yes it would have been easier. But i some what get it.

OpenStudy (anonymous):

smwhat is better than nothing